OpenStudy (anonymous):

lve: -1 over 2 x + 1 = −x + 8

OpenStudy (anonymous):

-1/(2x+1) or (-1/2x)+1?

OpenStudy (anonymous):

7 14 18 these are the only possible awnser

OpenStudy (anonymous):

and 1 over 2

OpenStudy (anonymous):

No post the proper question! There's a difference use parenthesis!

OpenStudy (anonymous):

its multiply choice... so i t has to be one of those.

OpenStudy (anonymous):

1 =-x+8 ---- 2x+1 1=(2x+1)(-x+8) 1=-2x^2+16x-x+8 1=-2x^2+15x+8 1-8=-2x^2+15x -7=-2x^2+15x -2x^2+15x=-7 -2x^2+15x+7=0 please do check my signs

OpenStudy (anonymous):

i still have no awnser...

OpenStudy (anonymous):

-1/(2x+1) = -x+8 -1 = -x+8(2x+1) -1 = -2x^2-x+16x+8 -1+8 = 2x^2+15x 7 = x(2x+15) 7/x = 2x+15 (7-15)/x = 2x -8/x = 2x -8 = 2x^2 x^2 = -8/2 x = root of -4 x = -2 but i am not sure about this

OpenStudy (anonymous):

-1 over 2 x + 1 = −x + 8 No wait i think he meant: (-1/2)x+1 = -x+8 (1/2)x+1 = 8 x+2= 16 x = 14

OpenStudy (anonymous):

THANKYOU

OpenStudy (anonymous):

so from now on, use parenthesis!